Write an equation of the given point and perpendicular to the given line.
(5,-6);6x+5y=8
The equation of the line is y=?

Answer by mananth(16946)

6x+5y=8
Find the slope of this line
5y=-6x-8
Divide by 5
y=-1 1/5x-1.6
Compare this equation with y=mx+b
slope m =-1.20

The slope of a line perpendicular to the above line will be the negative reciprocal 0.83
m1*m2=-1
The slope of the required line will be 0.83

m= 0.83 ,point (5,-6)
Find b by plugging the values of m & the point in
y=mx+b
-6=4.17+b
b=-10.17
m=0.83
The required equation is y=0.83x-10.17